# Motion Avatar: Generate Human and Animal Avatars with Arbitrary Motion

### v1

@@ 1 Introduction
Recent advancements in text-to-motion generation [Guo et al.(2022a)Guo, Zou, Zuo, Wang, Ji, Li, and Cheng, Tevet et al.(2022b)Tevet, Raab, Gordon, Shafir, Cohen-or, and Bermano, Chen et al.(2023b)Chen, Jiang, Liu, Huang, Fu, Chen, and Yu, Zhang et al.(2024b)Zhang, Liu, Reid, Hartley, Zhuang, and Tang, Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng], employing both diffusion [Ho et al.(2020)Ho, Jain, and Abbeel, Song et al.(2020)Song, Meng, and Ermon, Dhariwal and Nichol(2021), Nichol et al.(2021)Nichol, Dhariwal, Ramesh, Shyam, Mishkin, McGrew, Sutskever, and Chen, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer] and autoregressive [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin, Radford et al.(2018)Radford, Narasimhan, Salimans, Sutskever, et al.] models, have demonstrated significant promise. These advancements represent a notable step forward in generating motion sequences directly from textual descriptions, which brings a new approach to dynamic motion avatar generation. Meanwhile, significant advancements in 3D mesh reconstruction and generation [Qian et al.(2023)Qian, Mai, Hamdi, Ren, Siarohin, Li, Lee, Skorokhodov, Wonka, Tulyakov, et al., Liu et al.(2023b)Liu, Wu, Van Hoorick, Tokmakov, Zakharov, and Vondrick, Liu et al.(2024b)Liu, Xu, Jin, Chen, Varma T, Xu, and Su, Hong et al.(2023)Hong, Zhang, Gu, Bi, Zhou, Liu, Liu, Sunkavalli, Bui, and Tan, Tochilkin et al.(2024)Tochilkin, Pankratz, Liu, Huang, Letts, Li, Liang, Laforte, Jampani, and Cao], particularly through approach of image-to-3D, have demonstrated considerable promise and dominance in the field of 3D object reconstruction. These advancements hold great potential for real-world applications. However, existing efforts predominantly focus on either creating the 3D avatar mesh independently or generating motion sequences separately. Yet, effectively integrating these two components remains a persistent challenge.
@@ 2 Related Works
Creating motion is a vital aspect of computer vision, playing a crucial role in diverse applications such as 3D video animation, computer games, and robot manipulation. A prevalent approach to generating motion, known as the Text-to-Motion task, revolves around establishing a shared latent space for both textual descriptions and motion data. Among these applications, human motion generation has garnered significant attention and exploration. MotionCLIP [Tevet et al.(2022a)Tevet, Gordon, Hertz, Bermano, and Cohen-Or] leverages Transformer-based [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] Autoencoders to reconstruct motion sequences while ensuring alignment with corresponding textual labels within the CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] space. This alignment effectively integrates semantic knowledge from CLIP into the manifold of human motion. TEMOS [Petrovich et al.(2022)Petrovich, Black, and Varol] and T2M [Guo et al.(2022a)Guo, Zou, Zuo, Wang, Ji, Li, and Cheng] merge a Transformer-based Variational Autoencoder (VAE) [Kingma and Welling(2013)] with a text encoder to produce distribution parameters that operate within the latent space of the VAE. AttT2M [Zhong et al.(2023)Zhong, Hu, Zhang, and Xia] and TM2D [Gong et al.(2023)Gong, Lian, Chang, Guo, Jiang, Zuo, Mi, and Wang] integrate a body-part spatio-temporal encoder into VQ-VAE to facilitate improved learning of a discrete latent space with heightened expressiveness. MotionDiffuse [Zhang et al.(2024a)Zhang, Cai, Pan, Hong, Guo, Yang, and Liu] presents a pioneering framework for generating motion guided by text, leveraging diffusion models [Ho et al.(2020)Ho, Jain, and Abbeel, Song et al.(2020)Song, Meng, and Ermon, Dhariwal and Nichol(2021), Nichol et al.(2021)Nichol, Dhariwal, Ramesh, Shyam, Mishkin, McGrew, Sutskever, and Chen, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer]. It demonstrates various appealing traits such as probabilistic mapping, lifelike synthesis, and multi-tiered manipulation. Conversely, MDM [Tevet et al.(2022b)Tevet, Raab, Gordon, Shafir, Cohen-or, and Bermano] employs a Transformer-based diffusion model devoid of classifiers to forecast samples, rather than noise, at each step. In contrast, MLD [Chen et al.(2023b)Chen, Jiang, Liu, Huang, Fu, Chen, and Yu] conducts diffusion within latent motion space, eschewing the utilization of a diffusion model to forge links between raw motion sequences and conditional inputs. Motion Mamba [Zhang et al.(2024b)Zhang, Liu, Reid, Hartley, Zhuang, and Tang] proposed an approach which leveraged state space models (SSMs) [Gu and Dao(2023)] and diffusion to design an efficient architecture for long-sequence human motion generation. MoMask [Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng] proposed a novel masked modeling framework, designed a hierarchical quantization scheme architecture, and achieved superior results on text-driven 3D human motion generation.
@@ 3 Datasets
For tuning and evaluating our LLM planner, we’ve created the avatar Q&A dataset. This dataset utilizes animal and motion categories from Zoo-300K as input and is generated by LLaMA-70B [Touvron et al.(2023)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.]. It serves as a corpus for adjusting our LLM planner, with the instruction attribute used as prompts and the output attribute as the desired output. The dataset contains 1756 items, and Listing LABEL:lst:qa provides several examples. The dataset will assess whether the LLM planner can accurately identify animal and motion categories from natural language descriptions provided by users.
@@ 5.2 Motion Evaluation
Since human motion generation has already been evaluated in the MoMask [Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng], our focus shifts to comprehensively assessing animal motion generation utilizing the newly introduced Zoo-300K dataset. We will include a detailed quantitative evaluation in the upcoming revision of this paper. For qualitative assessment, the visualization in fig. 3 illustrates our model’s capability to seamlessly generate animal motion based on text conditions, showcasing potential applications in fields like computer gaming and film making. These findings highlight the versatility and effectiveness of our approach in diverse scenarios.
@@ 6 Discussion and Conclusion
Appendix A User Study

### v2

# Zeyu Zhanghttps://steve-zeyu-zhang.github.io12†

@@ Abstract
Project Website: https://steve-zeyu-zhang.github.io/MotionAvatar
@@ 1 Introduction
Recent advancements in text-to-motion generation [Guo et al.(2022a)Guo, Zou, Zuo, Wang, Ji, Li, and Cheng, Tevet et al.(2022b)Tevet, Raab, Gordon, Shafir, Cohen-or, and Bermano, Chen et al.(2023b)Chen, Jiang, Liu, Huang, Fu, Chen, and Yu, Zhang et al.(2024c)Zhang, Liu, Reid, Hartley, Zhuang, and Tang, Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng], employing both diffusion [Ho et al.(2020)Ho, Jain, and Abbeel, Song et al.(2020)Song, Meng, and Ermon, Dhariwal and Nichol(2021), Nichol et al.(2021)Nichol, Dhariwal, Ramesh, Shyam, Mishkin, McGrew, Sutskever, and Chen, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer] and autoregressive [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin, Radford et al.(2018)Radford, Narasimhan, Salimans, Sutskever, et al.] models, have demonstrated significant promise. These advancements represent a notable step forward in generating motion sequences directly from textual descriptions, which brings a new approach to dynamic motion avatar generation. Meanwhile, significant advancements in 3D mesh reconstruction and generation [Qian et al.(2023)Qian, Mai, Hamdi, Ren, Siarohin, Li, Lee, Skorokhodov, Wonka, Tulyakov, et al., Liu et al.(2023b)Liu, Wu, Van Hoorick, Tokmakov, Zakharov, and Vondrick, Liu et al.(2024b)Liu, Xu, Jin, Chen, Varma T, Xu, and Su, Hong et al.(2023)Hong, Zhang, Gu, Bi, Zhou, Liu, Liu, Sunkavalli, Bui, and Tan, Tochilkin et al.(2024)Tochilkin, Pankratz, Liu, Huang, Letts, Li, Liang, Laforte, Jampani, and Cao], particularly through approach of image-to-3D, have demonstrated considerable promise and dominance in the field of 3D object reconstruction. These advancements hold great potential for real-world applications. However, existing efforts predominantly focus on either creating the 3D avatar mesh independently or generating motion sequences separately. Yet, effectively integrating these two components remains a persistent challenge.
@@ 2 Related Works
Creating motion is a vital aspect of computer vision, playing a crucial role in diverse applications such as 3D video animation, computer games, and robot manipulation. A prevalent approach to generating motion, known as the Text-to-Motion task, revolves around establishing a shared latent space for both textual descriptions and motion data. Among these applications, human motion generation has garnered significant attention and exploration. MotionCLIP [Tevet et al.(2022a)Tevet, Gordon, Hertz, Bermano, and Cohen-Or] leverages Transformer-based [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] Autoencoders to reconstruct motion sequences while ensuring alignment with corresponding textual labels within the CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] space. This alignment effectively integrates semantic knowledge from CLIP into the manifold of human motion. TEMOS [Petrovich et al.(2022)Petrovich, Black, and Varol] and T2M [Guo et al.(2022a)Guo, Zou, Zuo, Wang, Ji, Li, and Cheng] merge a Transformer-based Variational Autoencoder (VAE) [Kingma and Welling(2013)] with a text encoder to produce distribution parameters that operate within the latent space of the VAE. AttT2M [Zhong et al.(2023)Zhong, Hu, Zhang, and Xia] and TM2D [Gong et al.(2023)Gong, Lian, Chang, Guo, Jiang, Zuo, Mi, and Wang] integrate a body-part spatio-temporal encoder into VQ-VAE to facilitate improved learning of a discrete latent space with heightened expressiveness. MotionDiffuse [Zhang et al.(2024a)Zhang, Cai, Pan, Hong, Guo, Yang, and Liu] presents a pioneering framework for generating motion guided by text, leveraging diffusion models [Ho et al.(2020)Ho, Jain, and Abbeel, Song et al.(2020)Song, Meng, and Ermon, Dhariwal and Nichol(2021), Nichol et al.(2021)Nichol, Dhariwal, Ramesh, Shyam, Mishkin, McGrew, Sutskever, and Chen, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer]. It demonstrates various appealing traits such as probabilistic mapping, lifelike synthesis, and multi-tiered manipulation. Conversely, MDM [Tevet et al.(2022b)Tevet, Raab, Gordon, Shafir, Cohen-or, and Bermano] employs a Transformer-based diffusion model devoid of classifiers to forecast samples, rather than noise, at each step. In contrast, MLD [Chen et al.(2023b)Chen, Jiang, Liu, Huang, Fu, Chen, and Yu] conducts diffusion within latent motion space, eschewing the utilization of a diffusion model to forge links between raw motion sequences and conditional inputs. Motion Mamba [Zhang et al.(2024c)Zhang, Liu, Reid, Hartley, Zhuang, and Tang] and InfiniMotion [Zhang et al.(2024b)Zhang, Liu, Chen, Chen, Reid, Hartley, Zhuang, and Tang] proposed approaches which leveraged state space models (SSMs) [Gu and Dao(2023)] and diffusion to design an efficient architecture for long-sequence human motion generation. MoMask [Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng] proposed a novel masked modeling framework, designed a hierarchical quantization scheme architecture, and achieved superior results on text-driven 3D human motion generation.
@@ 3 Datasets
For tuning and evaluating our LLM planner, we’ve created the avatar Q&A dataset. This dataset utilizes animal and motion categories from Zoo-300K as input and is generated by LLaMA-70B [Touvron et al.(2023)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.]. It serves as a corpus for adjusting our LLM planner, with the instruction attribute used as prompts and the output attribute as the desired output. The dataset contains 1756 items, and Listing LABEL:lst:qa in Appendix C provides several examples. The dataset will assess whether the LLM planner can accurately identify animal and motion categories from natural language descriptions provided by users.
@@ 5.2 Motion Evaluation
Since human motion generation has already been evaluated in the MoMask [Guo et al.(2023)Guo, Mu, Javed, Wang, and Cheng], our focus shifts to comprehensively assessing animal motion generation utilizing the newly introduced Zoo-300K dataset. For quantitative results, please refer to the table in Appendix A: Animal Motion Quantative Evaluation and Appendix B: User Study in supplementary. For qualitative assessment, the visualization in fig. 3 illustrates our model’s capability to seamlessly generate animal motion based on text conditions, showcasing potential applications in fields like computer gaming and film making. These findings highlight the versatility and effectiveness of our approach in diverse scenarios.
@@ 6 Discussion and Conclusion
Appendix A Animal Motion Quantative Evaluation
We evaluated our Motion Avatar on the Zoo-300K dataset, and the results are presented in table 2. The comprehensive evaluation demonstrated that our method achieves high-quality animal motion generation. The results indicate that our approach produces highly realistic and promising animations, highlighting the effectiveness and potential of our technique in generating detailed and lifelike animal motions.
| Method | R-Prec Top 1 | R-Prec Top 2 | R-Prec Top 3 | FID | MultiModal-Dist | Diversity |
| Anaconda (GT) | ||||||
| Anaconda (Ours) | ||||||
| Ant (GT) | ||||||
| Ant (Ours) | ||||||
| Bat (GT) | ||||||
| Bat (Ours) | ||||||
| Bear (GT) | ||||||
| Bear (Ours) | ||||||
| Bird (GT) | ||||||
| Bird (Ours) | ||||||
| Buffalo (GT) | ||||||
| Buffalo (Ours) | ||||||
| Buzzard (GT) | ||||||
| Buzzard (Ours) | ||||||
| Camel (GT) | ||||||
| Camel (Ours) | ||||||
| Cat (GT) | ||||||
| Cat (Ours) | ||||||
| Centipede (GT) | ||||||
| Centipede (Ours) | ||||||
| Chicken (GT) | ||||||
| Chicken (Ours) | ||||||
| Cobra (GT) | ||||||
| Cobra (Ours) | ||||||
| Komodo (GT) | ||||||
| Komodo (Ours) | ||||||
| Coyote (GT) | ||||||
| Coyote (Ours) | ||||||
| Crab (GT) | ||||||
| Crab (Ours) | ||||||
| Cricket (GT) | ||||||
| Cricket (Ours) | ||||||
| Crocodile (GT) | ||||||
| Crocodile (Ours) | ||||||
| Crow (GT) | ||||||
| Crow (Ours) | ||||||
| Deer (GT) | ||||||
| Deer (Ours) | ||||||
| Dog (GT) | ||||||
| Dog (Ours) | ||||||
| Eagle (GT) | ||||||
| Eagle (Ours) | ||||||
| Elephant (GT) | ||||||
| Elephant (Ours) | ||||||
| Fire Ant (GT) | ||||||
| Fire Ant (Ours) | ||||||
| Flamingo (GT) | ||||||
| Flamingo (Ours) | ||||||
| Fox (GT) | ||||||
| Fox (Ours) |
| Method | R-Prec Top 1 | R-Prec Top 2 | R-Prec Top 3 | FID | MultiModal-Dist | Diversity |
| Gazelle (GT) | ||||||
| Gazelle (Ours) | ||||||
| Giant Bee (GT) | ||||||
| Giant Bee (Ours) | ||||||
| Goat (GT) | ||||||
| Goat (Ours) | ||||||
| Hamster (GT) | ||||||
| Hamster (Ours) | ||||||
| Hermit Crab (GT) | ||||||
| Hermit Crab (Ours) | ||||||
| Hippopotamus (GT) | ||||||
| Hippopotamus (Ours) | ||||||
| Horse (GT) | ||||||
| Horse (Ours) | ||||||
| Hound (GT) | ||||||
| Hound (Ours) | ||||||
| Isopetra (GT) | ||||||
| Isoptera (Ours) | ||||||
| Jaguar (GT) | ||||||
| Jaguar (Ours) | ||||||
| Leopard (GT) | ||||||
| Leopard (Ours) | ||||||
| Lion (GT) | ||||||
| Lion (Ours) | ||||||
| Lynx (GT) | ||||||
| Lynx (Ours) | ||||||
| Mammoth (GT) | ||||||
| Mammoth (Ours) | ||||||
| Monkey (GT) | ||||||
| Monkey (Ours) | ||||||
| Ostrich (GT) | ||||||
| Ostrich (Ours) | ||||||
| Parrot (GT) | ||||||
| Parrot (Ours) | ||||||
| Pigeon (GT) | ||||||
| Pigeon (Ours) | ||||||
| Piranha (GT) | ||||||
| Piranha (Ours) | ||||||
| Polar Bear (GT) | ||||||
| Polar Bear (Ours) | ||||||
| Pteranodon (GT) | ||||||
| Pteranodon (Ours) | ||||||
| Puppy (GT) | ||||||
| Puppy (Ours) | ||||||
| Raptor (GT) | ||||||
| Raptor (Ours) | ||||||
| Rat (GT) | ||||||
| Rat (Ours) | ||||||
| Reindeer (GT) | ||||||
| Reindeer (Ours) |
| Method | R-Prec Top 1 | R-Prec Top 2 | R-Prec Top 3 | FID | MultiModal-Dist | Diversity |
| Rhino (GT) | ||||||
| Rhino (Ours) | ||||||
| Roach (GT) | ||||||
| Roach (Ours) | ||||||
| Sabre-toothed tiger (GT) | ||||||
| Sabre-toothed Tiger (Ours) | ||||||
| Sand mouse (GT) | ||||||
| Sand Mouse (Ours) | ||||||
| Scorpion (GT) | ||||||
| Scorpion (Ours) | ||||||
| Shark (GT) | ||||||
| Shark (Ours) | ||||||
| Skunk (GT) | ||||||
| Skunk (Ours) | ||||||
| Spider (GT) | ||||||
| Spider (Ours) | ||||||
| Stegosaurus (GT) | ||||||
| Stegosaurus (Ours) | ||||||
| T-Rex (GT) | ||||||
| T-Rex (Ours) | ||||||
| Tricera (GT) | ||||||
| Tricera (Ours) | ||||||
| Toucan (GT) | ||||||
| Toucan (Ours) | ||||||
| Turtle (GT) | ||||||
| Turtle (Ours) | ||||||
| Tyrannosaurus Rex (GT) | ||||||
| Tyrannosaurus Rex (Ours) | ||||||
| Wyvern (GT) | ||||||
| Wyvern (Ours) | ||||||
| Average (GT) | ||||||
| Average (Ours) |
Appendix B User Study